OpenStudy (anonymous):

[Part 1] Show all work in dividing the fraction of 3 k squared plus k minus 2 over 4 k minus 2 divided by the fraction of k squared plus 3 k plus 2 all over 2 k squared plus 5 k minus 3. [Part 2] Use complete sentences to explain your process.

OpenStudy (anonymous):

multiply by the reciprocal just like you would do with numbers \[\frac{3k^2+k+2}{4k^2-2}\times \frac{2k^2+5k-3}{k^2+3k+2}\]

OpenStudy (anonymous):

then factor and cancel like before

OpenStudy (anonymous):

\[\frac{3k^2+k+2}{4k-2}\times \frac{2k^2+5k-3}{k^2+3k+2}\]

OpenStudy (anonymous):

@satellite73 are you a teacher? student?

OpenStudy (anonymous):

typo in my first one @Trexy will show you how to factor

OpenStudy (anonymous):

i am a car

OpenStudy (anonymous):

@HayleyWhite17 you got the factoring part?

OpenStudy (anonymous):

No really, Im asking cuz their is a diffrent between a teacher and a student, you know what i mean..@satellite73

OpenStudy (anonymous):

No not quite.

OpenStudy (anonymous):

\(x\) should be \(k\)

OpenStudy (anonymous):

dang lots of typos but you can fix it

OpenStudy (anonymous):

\[=\frac{(k+1)(3k-2)(k+3)(2k-1)}{2(2k-1)(k+2)(k+1)}\]

OpenStudy (anonymous):

No I get what it is THANKS!